OpenStudy (kweenlexis):

5 question test...Help please will metall and fan!! question in comment because i have to personally type it in.

OpenStudy (kweenlexis):

1. What is tthe LCD of \[\frac{ 1 }{ 3 }\] and \[\frac{ 3}{ 4 }\] A.\[3\] B.\[4\] C.\[7\] D.\[12\] I am going to go with D am i correct?

OpenStudy (anonymous):

A

OpenStudy (kweenlexis):

What is the LCD of \[\frac{ 5 }{ 12 }\] and \[- \frac{ 9 }{ 16 }\]

OpenStudy (kweenlexis):

How did you get A??

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

im sorry your right it d i was looking at some thing else

OpenStudy (kweenlexis):

its okay thanks

OpenStudy (anonymous):

5/12,-9/16 lcd is 48